The cost of electricity required to deposit 1 gram of Mg = Rs 5 .determine how much will it cost to deposit 10 grams of aluminium? 

Grade:

2 Answers

Arun
4 years ago
1 Gram of Mg means 0.04 Moles i.e 2.5 × 10^22 Atoms. Whose cost is 5 Rupees. 10g of Al contains 2.2 ×10^23 Atoms. Therefore the cost is 5/(2.5× 10^22) * 2.2 *10^23 Which is 44
Khimraj
4 years ago
m=eQ for Mg
m=EQ/F
Q(RS 5)=1F/12
m=eQ for Al
m=EQ/F
Q = 10F/9 = 10F/9X1/12X12/1=66.66
Hope it clears...................................................